# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ C.1 Additional Verification Experiments
Model indistinguishability: For FineTune, we can observe that even though the accuracy results are similar to the retrained model, the distance metric is very large (e.g. 2.21 on CIFAR-10 and 2.36 on CelebA). For NegGrad, we observe that the accuracy on forgetting dataset $D_{f}$ and remaining dataset $D_{r}$ would rapidly decrease simultaneously, but even in this case, the distances are is still lower than that of FineTune for all three datasets. This demonstrates that a small difference in accuracy does not indicate the model indistinguishability. Therefore, we use the $L_{2}$ norm and correlation metric as the primary measure of forgetting completeness, which are commonly employed in studies related to influence function [35, 42]. These studies serve as inspiration for prior theoretical certified unlearning works [2, 3, 4] that we compare. Measuring forgetting through $L_{2}$ norm is one of the most stringent metrics and is also used in unlearning research, e.g., [15, 18], but it necessitates retraining a model, which is impractical for real-world applications. Therefore, exploring more effective and efficient metrics to evaluate model indistinguishability is an interesting and important topic in the future.
@@
C.2 Additional Application Experiments
@@
Furthermore, we evaluate the performance on different datasets and models in real-world application. Our evaluations are conduct from two perspectives: runtime and utility. Specifically, run-time focuses on the time spent the precomputing unlearning statistics and the speedup of the unlearning algorithm compared to the retraining algorithm. Moreover, we evaluate the utility by the test accuracy of the unlearned model to ensure that generalization performance is not compromised.
@@
Configurations: We assess the performance on various datasets and models, as presented in Table 6-Table 7. The values in parentheses indicate the difference in test accuracy compared to the retrained model. (1) We train a LR and simple CNN on MNIST with 1,000 data points with 20% data to be forgotten, which have setups identical to the aforementioned verification experiments I. We further simulate on FMNIST with 4,000 data using CNN and LeNet with total 61,706 parameters. The training is conducted for 30 epochs with stepsize of 0.5 and a batch size of 256. (2) Moreover, we assessed our method on Resnet-18 [47] trained on the CIFAR10 dataset for image classification with 50,000 data points, CelebA dataset for gender prediction with 10,000 data points, and LFWPeople dataset for face recognition with 984 data points. The hyperparameter configurations align with those detailed in Table 3-5. In these circumstances, we have not evaluated NS and IJ due to out-of-memory.
@@
Application experiment results show that our proposed method is computationally efficient and performs well in realistic environments, outperforming benchmarks of certified unlearning regarding all metrics by a significant margin, as demonstrated in Table 6. It achieves minimal computational and storage overhead while maintaining high performance in both convex and non-convex settings. We further investigate data removal with large-scale datasets using ResNet-18, as shown in Table 7. Our proposed Hessian-free method demonstrates significant potential for over-parameterized models, as previous second-order certified unlearning works are unable to compute the full Hessian matrix in such cases. Notably, for unlearning runtime, by precomputing statistical data before unlearning requests arrive, we only need to execute simple vector additions, achieving millisecond-level efficiency (0.6 ms) with minimal performance degradation in test accuracy.
Finally, although our method has lower computation and storage complexity compared to previous works, it still requires $\mathcal{O}(nTd)$ precomputation and $\mathcal{O}(nd)$ storage which is expensive when facing both larger-scale models and datasets. We summarize the limitations in Subsection E.2, and provide possible schemes to address these problems.
Table 6: Application experiments II. An online scenario on small-scale datasets and models, where the unlearning requests involve only forgetting a single data point with each execution. We compute the total computational/storage cost for all samples and tested the accuracy of final unlearned model.
Method Dataset Model Unlearning Computaion PreComputaion Storage Dis. Test Accuracy (%)
Runtime (Sec) Speedup Runtime (Sec) (GB) Unlearned model
NS MNIST Logistic 5.14×$\times 10^{2}$ 2.39$\times$ 2.57×$\times 10^{3}$ 0.23 0.13 87.50 (-0.75)
CNN 5.82×$\times 10^{3}$ 0.05$\times$ 2.91×$\times 10^{4}$ 1.78 1.78 83.50 (-10.25)
FMNIST CNN 2.31×$\times 10^{4}$ 0.54$\times$ 1.16×$\times 10^{5}$ 1.78 2.41 57.69 (-22.25)
LeNet 8.53×$\times 10^{4}$ 0.15$\times$ 4.27×$\times 10^{5}$ 14.18 1.10 62.00 (-18.50)
IJ MNIST Logistic 1.84×$\times 10^{0}$ 665$\times$ 2.57×$\times 10^{3}$ 0.25 0.13 87.50 (-0.75)
CNN 6.81×$\times 10^{0}$ 45$\times$ 2.91×$\times 10^{4}$ 1.78 3.12 82.75 (-11.00)
FMNIST CNN 2.95×$\times 10^{1}$ 424$\times$ 1.16×$\times 10^{5}$ 1.78 0.94 75.69 (-4.25)
LeNet 3.24×$\times 10^{1}$ 402$\times$ 4.27×$\times 10^{5}$ 14.18 0.57 76.75 (-3.75)
Ours MNIST Logistic 2.74×$\times 10^{-3}$ 448,060$\times$ 1.95×$\times 10^{1}$ 0.03 0.12 87.75 (-0.50)
CNN 3.23×$\times 10^{-2}$ 9,940$\times$ 5.34×$\times 10^{2}$ 0.08 0.68 91.50 (-2.25)
FMNIST CNN 1.27×$\times 10^{-1}$ 98,468$\times$ 3.66×$\times 10^{3}$ 0.32 0.74 77.85 (-2.09)
LeNet 1.63×$\times 10^{-1}$ 79,828$\times$ 4.43×$\times 10^{3}$ 0.48 0.53 78.63 (-1.87)
@@
Table 7: Application experiments III. An online scenario on the Resnet-18 with large-scale datasets. We compute the computational and cost overhead of per forgetting sample.
@@
Method Dataset Unlearning Computaion PreComputaion Storage Dis. Test Accuracy (%)
(per sample) (per sample) (per sample)
Runtime (Sec) Speedup Runtime (Sec) (GB) Unlearned model
Ours CIFAR10 0.6×$\times 10^{-3}$ 781,083$\times$ 184.53 0.043 0.09 79.62 (-0.01)
LFW 1.0×$\times 10^{-3}$ 77,430 $\times$ 84.99 0.043 0.31 71.92 (-8.97)
CelebA 0.6×$\times 10^{-3}$ 146,200$\times$ 31.90 0.042 0.29 95.57 (-0.15)
Appendix D Ablation Studies
@@
Here, we investigate the sensitivity of the proposed approach to hyperparameters. Specifically, we evaluated the effects of different step sizes, epochs, decay rates of step size, and dataset sizes on proposed approach. We conducted simulations in both convex and non-convex scenarios, using Logistic Regression for the convex case and CNN for the non-convex case, respectively.
@@
Takeaway. Let’s first provide a brief summary for the main results of ablation studies.
  • Impact of Step Sizes. Firstly, a smaller step size enables our model to be closer to the retrained model. This is because Theorem 4 demonstrates that the approximation error mainly consists of two components: (1) the step size $\eta_{e,b}$ and (2) the maximum eigenvalue $\lambda_{1}$ of $\frac{1}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$. Therefore, a smaller stepsize $\eta_{e,b}$ constrain the maximum eigenvalue of the matrix $\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, leading to a reduction in the error by decreasing the value of $\rho$. Moreover, with a small step size, increasing epochs does not lead to significant approximation errors; instead, it exhibits better performance under the distance metric compared to a large stepsize.
  • Impact of Epochs. Secondly, when the step size is sufficiently small, the approximation error is insensitive and exhibits slow growth with respect to the number of epochs, i.e., excessive iterations in this scenario do not lead to significant errors.
  • Impact of Stepsize Decay Rates. Thirdly, our results indicate that as the decay rate $q$ decreases, the approximation error diminishes. Specifically, The stepsize decay strategy introduces a bound on approximation error, preventing a continuous increase in error as the size of epochs grows.
  • Impact of Dataset Sizes. Finally, our numerical findings show that the approximation error reduces with larger dataset sizes. Intuitively, as the dataset size increases, the impact of individual data points on model updates diminishes during training. This is typically reflected in the training process, where with an increase in the dataset size, each batch update is less likely to frequently sample a specific data point. Consequently, the impact of any individual data point on updates between the retrained original models diminishes, leading to a more accurate approximation.
D.1 Impact of Step Sizes
@@
In this section, we investigated (i) the impact of step sizes and (ii) the impact of small step size with more epochs. Additionally, we discussed the connection between sharpness in learning generalization studies and our unlearning method.
Configurations: (i) We investigate the impact of step sizes {0.001, 0.005, 0.01, 0.05, 0.1, 0.5} on distance metrics in Table 9 and Table 9. (ii) Furthermore, in Figure 3, to investigate the effect of epochs under small step sizes, LR was trained for 750 epochs and 2000 epochs with a step size of 0.001 (initially 0.05 with 50 epochs). For CNN, training for 50 epochs and 500 epochs with a step size of 0.025 (originally 0.05 with 20 epochs) was performed. Additionally, when the step size is very small (0.001) in LR, the stepsize decay strategy causes premature decay to zero before the model convergence. We thus increased the stepsize decay rate from 0.995 to 1. Subsequent experiments will demonstrate that this leads to an increase in approximation error. Finally, we keep the remaining hyperparameters consistent with verification experiment I.
@@
Impact of step sizes. As shown in Table 9 and Table 9, a smaller step size enables our model to be closer to the retrained model. Specifically, as Theorem 4 demonstrates that, in this process, the error arises from the scaled remainder term, i.e., the term $o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})$ is scaled by the subsequent $\mathbf{M}_{e,b(u)}$, where $\mathbf{M}_{e,b(u)}=\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\prod^{E}_{k=e}\prod_{b=b(u)+1}^{B-1}\Big{(}\mathbf{I}-\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b}\Big{)}$. Therefore, the approximation error mainly consists of two components: (1) the step size $\eta_{e,b}$ and (2) the maximum eigenvalue $\lambda_{1}$ of $\frac{1}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$. Therefore, a smaller stepsize $\eta$ constrain the maximum eigenvalue of the matrix $\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, thereby reducing the spectral radius $\rho$ of the matrix $\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, and finally reducing approximation error.
@@
Table 8: Impact of Step Sizes on Convex Setting. Stepsize $\eta=0.05$ represent hyperparameter choices of previous experiments for LR in our paper. It can be observed that as the step size decreases, our method’s unlearned model becomes closer to the retrained model. (Seed 42)
@@
0.001 0.002 0.005 0.014 0.020
0.005 0.007 0.017 0.046 0.065
0.01 0.010 0.024 0.067 0.097
0.05 0.030 0.060 0.163 0.230
0.1 0.047 0.095 0.250 0.348
0.5 0.800 1.730 12.533 14.825
Table 9: Impact of Step Sizes on Non-Convex Setting. Stepsize $\eta=0.05$ represent hyperparameter choices of previous experiments for CNN in our paper. It can be observed that as the step size decreases, our method’s unlearned model becomes closer to the retrained model. (Seed 42)
@@
0.1 0.437 0.795 1.722 2.725
0.5 8E+7 3E+7 2E+8 7E+7
@@
Figure 3: Impact of Small Step Size with More Epochs. (a)(b) correspond to LR and CNN with 20% of the data to be forgotten, respectively. The red bar in the left represents the distance metric when using a larger step size; the blue bar in the middle represents the distance metric with a small stepsize during the large epochs; and the right light blue bar in the right represent the distance metric with a small stepsize during the small epochs. It can be observed that smaller step sizes lead to smaller approximation errors. Furthermore, even when training more epochs with smaller step sizes, the approximation error remains lower compared to training with larger learning rates. (Seed 42)
Furthermore, if the step size is extremely large, it will result in a large spectral radius $\rho$ of the matrix $\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, causing exponential growth in errors, as demonstrated in Table 9 and Table 9 for an initial step size of $\eta=0.5$. In this scenario, our method leads to a complete breakdown of the model, rendering it unusable. However, it is noteworthy that theoretical predictions [51] and empirical validation [52] suggest that, across all datasets and models, successful training occurs only when optimization enters a stable region of parameter space where $\lambda_{1}\cdot\eta\leq 2$. Therefore, achieving good optimization results typically does not require a strategy of using aggressive step sizes.
The connection between learning and unlearning. There is a connection between the sharpness of learning generalization studies and the unlearning method that we propose. Specifically, these studies aim to explore the sharpness of the Hessian of loss function $\frac{1}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, and the term ’sharpness’ refers to the maximum eigenvalue of the Hessian, denoted as $\lambda_{1}$ in some studies, such as [52]. A flat loss space is widely recognized as beneficial for gradient descent optimization, leading to improved convergence. Importantly, [52] demonstrates the central role that $\lambda_{1}$ plays in neural network optimization, emphasizing that maintaining a sufficiently small $\lambda_{1}$ during optimization is a necessary condition for successful training (without causing divergence) at large step sizes. Models which train successfully enter a region where $\lambda_{1}\cdot\eta\leq 2$ mid training (or fluctuate just above this bound).
In our work, we leverage the second-order information from the loss Hessian at each iteration during the training process to compute the approximators and achieve forgetting. As indicated in Theorem 4, our analysis indicates that a small $\lambda_{1}$ of the Hessian is beneficial for reducing the approximation error. Therefore, effectively reducing $\lambda_{1}$ during optimization, so that the training enters a flatter region, will not only improve the generalization ability of the learning stage model but also reduce the approximation error in the unlearning stage. As suggested in [52], strategies like stepsize warm-up or initialization strategies for architectures can be employed to stabilize learning and potentially reduce unlearning approximation errors by decreasing $\lambda_{1}$, enabling training at larger step sizes.
Impact of small step size with more epochs. Finally, with a very small step size during model training, we observed a low approximation error. As shown in Figure 3, even as in the case with large epoch sizes, the approximation error remains lower than in the case with a large step size of 0.05. This suggests that a smaller step size reduces the error. It is noted that a smaller step size requires more iterations to converge, resulting in increased computational cost. Therefore, trading an appropriate step size is essential based on different scenarios.
D.2 Impact of Epochs
Our aforementioned experiments indicate that the increase in epochs does not become a major factor affecting approximation errors when applying small step sizes, i.e., with LR at a step size of 0.001 and CNN at a step size of 0.025, the growth of epochs does not significantly impact the approximation error. Conversely, in the case of small step sizes, even as epochs extend to a point where the model approaches overfitting, the error remains lower than scenarios with larger step sizes.
Configurations: Furthermore, we explore the impact of the epoch with a fixed and larger step size, such as the one $\eta=0.05$ we selected in our study. In this scenario, we evaluate the impact of the epochs on both the distance and accuracy metrics. Specifically, we investigated the distance under the epoch set {10, 25, 50, 75, 100} and provided the accuracy on 100 epoch after forgetting 20% data. In addition, the selection of other hyperparameters remains consistent with the ones used previously.

### v4

@@ D.2 Additional Verification Experiments
Model indistinguishability: For FineTune, we can observe that even though the accuracy results are similar to the retrained model, the distance metric is very large (e.g. 2.21 on CIFAR-10 and 2.36 on CelebA). For NegGrad, we observe that the accuracy on forgetting dataset $D_{f}$ and remaining dataset $D_{r}$ would rapidly decrease simultaneously, but even in this case, the distances are is still lower than that of FineTune for all three datasets. This demonstrates that a small difference in accuracy does not indicate the model indistinguishability. Therefore, we use the $L^{2}$ norm and correlation metric as the primary measure of forgetting completeness, which are commonly employed in studies related to influence function (Basu et al. (2021; 2020)). These studies serve as inspiration for prior certified unlearning papers (Sekhari et al. (2021); Suriyakumar et al. (2022); Liu et al. (2023)) that we compare. Measuring forgetting through $L^{2}$ norm is also used in unlearning research, e.g., Wu et al. (2020); Izzo et al. (2021), but it necessitates retraining a model, which is impractical for real-world applications.
@@
D.3 Additional Application Experiments
@@
Furthermore, we evaluate the performance on different datasets and models in real-world applications. Our evaluations are conducted from two perspectives: runtime and utility. Specifically, run-time focuses on the time spent precomputing unlearning statistics and the speedup of the unlearning algorithm compared to the retraining algorithm. Moreover, we evaluate the utility by the test accuracy of the unlearned model to ensure that generalization performance is not compromised. Finally, we record the area under the curve (AUC) for the MIA-L attack.
@@
Configurations: We assess the performance on various datasets and models, as presented in Table 6. The values in parentheses indicate the difference in test accuracy compared to the retrained model. (1) We train a LR and simple CNN on MNIST with 1,000 training data 20% data points to be forgotten, which have setups identical to the aforementioned verification experiments I. We further evaluate on FMNIST with 4,000 training data and 20% data points to be forgotten using CNN and LeNet with a total of 61,706 parameters. The training is conducted for 30 epochs with a stepsize of 0.5 and a batch size of 256. (2) Moreover, we assessed our method on Resnet-18 (He et al. (2016)) trained on the CIFAR10 dataset for image classification with 50,000 data points, CelebA dataset for gender prediction with 10,000 data points, and LFWPeople dataset for face recognition with 984 data points. we uniformly add noise with a standard deviation of 0.01. The configurations align with those detailed in Table 3-5. In these circumstances, we have not evaluated NS and IJ due to out-of-memory.
@@
Application experiment results show that our proposed method is computationally efficient and performs well in realistic environments, outperforming benchmarks of certified unlearning regarding all metrics by a significant margin, as demonstrated in Table 6. It achieves minimal computational and storage overhead while maintaining high performance in both convex and non-convex settings. We further investigate data removal with large-scale datasets using ResNet-18, as shown in Table 6. Our proposed Hessian-free method demonstrates significant potential for over-parameterized models, as previous second-order certified unlearning works are unable to compute the full Hessian matrix in such cases. Notably, for unlearning runtime, by precomputing statistical data before unlearning requests arrive, we only need to execute simple vector additions, achieving millisecond-level efficiency (0.6 ms) with minimal performance degradation in test accuracy. Finally, although our method has lower computation and storage complexity compared to previous works, it still requires $\mathcal{O}(n^{2}d)$ precomputation and $\mathcal{O}(nd)$ storage which is expensive when facing both larger-scale datasets. We summarize the limitations in Subsection B.1, and provide possible schemes to address these problems.
@@
Table 6: Application experiments II. We conduct two types of comparisons: (1) on small-scale models and datasets with previous second-order algorithms, where we compute the total computational/storage cost for all samples; and (2) on large-scale models and datasets with fast unlearning algorithms, where we compute the computational/cost overhead of per-sample forgetting.
@@
Method Dataset Model Unlearning Computaion PreComputaion Storage MIA-L Distance Test Accuracy (%)
Runtime (Sec) Speedup Runtime (Sec) (GB) AUC Unlearned model
NS MNIST LR 5.09×$\times 10^{2}$ 1.34$\times$ 2.55×$\times 10^{3}$ 0.23 0.51 0.18 86.25 (-1.50)
CNN 5.81×$\times 10^{3}$ 0.12$\times$ 2.91×$\times 10^{4}$ 1.78 0.51 1.78 83.50 (-10.25)
FMNIST CNN 2.31×$\times 10^{4}$ 0.54$\times$ 1.16×$\times 10^{5}$ 1.78 0.54 2.41 57.69 (-22.25)
LeNet 8.53×$\times 10^{4}$ 0.15$\times$ 4.27×$\times 10^{5}$ 14.18 0.54 1.10 62.00 (-18.50)
IJ MNIST LR 0.23×$\times 10^{2}$ 29.64$\times$ 2.55×$\times 10^{3}$ 0.23 0.52 0.18 86.25 (-1.50)
CNN 1.91×$\times 10^{2}$ 3.63$\times$ 2.91×$\times 10^{4}$ 1.78 0.51 3.12 82.75 (-11.00)
FMNIST CNN 2.95×$\times 10^{1}$ 424$\times$ 1.16×$\times 10^{5}$ 1.78 0.51 0.94 75.69 (-4.25)
LeNet 3.24×$\times 10^{1}$ 402$\times$ 4.27×$\times 10^{5}$ 14.18 0.54 0.57 76.75 (-3.75)
HF MNIST LR 0.0073 93,376$\times$ 2.20×$\times 10^{2}$ 0.03 0.51 0.15 87.50 (-0.25)
CNN 0.0268 25,821$\times$ 5.34×$\times 10^{2}$ 0.08 0.51 0.68 91.50 (-2.25)
FMNIST CNN 0.127 98,468$\times$ 3.66×$\times 10^{3}$ 0.32 0.51 0.74 77.85 (-2.09)
LeNet 0.163 79,828$\times$ 4.43×$\times 10^{3}$ 0.48 0.53 0.53 78.63 (-1.87)
Method Dataset Model Unlearning Computaion PreComputaion Storage MIA-L Distance Test Accuracy (%)
Runtime (Sec) Speedup Runtime (Sec) (GB) AUC Unlearned model
FineTune CIFAR10 106.76 4.39 $\times$ 0.56 2.21 79.34 (-0.29)
LFW Resnet18 15.60 4.96 $\times$ 0.53 0.52 76.02 (-4.87)
CelebA 35.09 2.50 $\times$ 0.56 2.36 95.78 (0.06)
NegGrad CIFAR10 0.56 836.88 $\times$ 0.51 0.11 40.61 (-39.03)
LFW Resnet18 0.46 168.33 $\times$ 0.46 0.48 63.01 (-17.88)
CelebA 0.96 91.38 $\times$ 0.50 1.13 37.95 (-57.77)
HF CIFAR10 0.0006 781,083$\times$ 184.53 0.043 0.51 0.06 79.62 (-0.01)
LFW Resnet18 0.001 77,430 $\times$ 84.99 0.043 0.56 0.31 71.92 (-8.97)
CelebA 0.0006 146,200$\times$ 31.90 0.042 0.54 0.29 95.57 (-0.15)
D.4 Additional Adversary Experiments
Membership Inference Attack against Learning (MIA-L) aims to determine whether a specific example is part of the training set of the learned model. Carlini et al. (2022) introduced the Likelihood Ratio Attack (LiRA), which leverages the model’s per-example difficulty scores and employs a well-calibrated Gaussian likelihood estimate to gauge the probability of a sample being part of training data. Finally, the adversary queries the confidence of the target model on a specific example and outputs a parametric likelihood-ratio test. Membership Inference Attack against Unearning (MIA-U) seeks to predict whether a target sample is part of the learned model’s training set but not in the corresponding unlearned (retrained) model’s training set. Chen et al. (2021) proposed MIA-U to attack Scratch methods, specifically the Retrain and SISA (Bourtoule et al. (2021)) algorithms, where the adversary jointly utilizes the posteriors of the two models as input to the attack model, either by concatenating them or by computing difference. Specifically, it includes five methods to construct features for the attack model: Direct difference (DirectDiff), Sorted difference (SortedDiff), Direct concatenate (DirectConcat), Sorted concatenate (SortedConcat), and Euclidean distance (EucDist).
Configuration: (1) We investigate the Area Under the Curve (AUC) of MIA-U at different deletion rates. (2) We report the Receiver Operating Characteristic (ROC) curve of MIA-U in large-scale experiments as well. (3) Finally, as a complement to the results of the DirectDiff method reported in the main text, we present the ROC of the remaining four feature construction methods in MIA-U. We maintain the other hyperparameters consistent with previous experiments.
Different deletion rates. As illustrated in Figure 5, a clear correlation exists between the deletion rate and the effectiveness of MIA-U attacks. Specifically, at a low deletion rate of 1%, the AUC of MIA-U attacks reaches 91% for our proposed method and 89% for the retrain method in the LR model, indicating that unlearning algorithms may increase the risk of privacy breaches. When the deletion rate exceeds 5%, the performance of MIA-U attacks approaches random guessing (AUC = 50%), suggesting that removing a larger amount of data between successive model releases makes it increasingly difficult for MIA-U attacks to accurately infer the deleted individual information.
Figure 5: Membership Inference Attack II. The AUC values of MIA-U at different deletion rates. The feature construction method employed in MIA-U is DirectDiff, with error bars representing 95% confidence intervals. The target model in the first plot is LR, while the second plot corresponds to a CNN. The findings show that lower deletion rates lead to privacy leakage, whereas higher deletion rates diminish MIA-U’s attack performance, bringing it closer to random guessing.
Figure 6: Membership Inference Attack III. ROC curves for MIA-U on ResNet18. The datasets used in the figure, from left to right, are CIFAR-10, LFW, and CelebA. For large-scale setups, our method (red line) performs consistently with the retraining algorithm (light red line) under MIA-U.
Different feature construction methods. As shown in Figure 7, we experiment with a well-generalized LR model and an overfitted CNN model, and the experimental results are consistent with the conclusion of Chen et al. (2021): sorting the posteriors improves attack performance; concatenation-based methods achieve higher attack performance on the overfitted CNN; and difference-based methods perform better on the well-generalized LR. In addition, we observe that perturbing the unlearning model can effectively defend the EucDist method. However, for other feature construction methods, adding noise to perturb the model after a certain level does not reduce the AUC of the attack model. This reveals a fundamental dilemma of most certified data removal processes developed so far: Although the goal is to achieve performance similar to a gold-standard retrained algorithm, the retrained model itself remains vulnerable to MIA-U, exposing the failure modes of certified unlearning. Nonetheless, as shown in Figure 5, removing at least 5% of the data when releasing the unlearned model can mitigate privacy leakage. However, we are not aware of any recent certified data removal mechanisms that can satisfy the deletion capacity exceeding 5% of the training set. Fortunately, one of the defense mechanisms proposed by Chen et al. (2021), which employs Differentially Private Stochastic Gradient Descent (DP-SGD), can effectively mitigate MIA-U but requires careful trade-off amongst privacy, unlearning efficiency, and model utility.
(a) SortedDif
(b) DirectConcat
(c) SortedConcat
(d) EucDist
Figure 7: Membership Inference Attack IV. ROC curve for different feature construction methods. The target model in the first row is overfitted CNN, whereas the second row is well-generalized LR.
Appendix E Sensitivity Analysis and Ablation Studies
@@
In this section, we investigate the effects of different step sizes, epochs, stepsize decay rates, and gradient clipping on the proposed approach. We conducted simulations in both convex and nonconvex scenarios, using LR for the convex case and 2-layer CNN for the non-convex case, respectively.
@@
Takeaway. Let’s first provide a brief summary of the main results of ablation studies.
(1) Impact of Step Sizes. Firstly, a smaller step size enables our model to be closer to the retrained model. However, when the step size exceeds a certain threshold, the bound will become meaningless and result in significant approximation errors. (2) Impact of Epochs. Secondly, when the step size is sufficiently small, the approximation error is insensitive and exhibits slow growth with respect to the number of epochs, i.e. excessive iterations in this scenario do not lead to significant errors. (3) Impact of Stepsize Decay Rates. Thirdly, our results indicate that as the decay rate $q$ decreases, the approximation error diminishes. Specifically, The stepsize decay strategy introduces a bound on approximation error, preventing a continuous increase in error as the size of epochs grows. (4) Impact of Gradient Clipping. To prevent gradient explosion, we introduced gradient clipping. We show that gradient clipping, which scales the gradient based on its norm, has a similar effect to using a smaller step size. Although, in theory, clipping the gradient may introduce errors, our experiments indicate that this impact is negligible when the threshold is set to a large value.
E.1 Impact of Step Sizes
@@
In this section, we investigated (i) the impact of step sizes and (ii) the impact of small stepsizes with more epochs. Additionally, we discussed the connection between learning generalization and our unlearning method.
Configurations: (i) We investigate the impact of step sizes {0.005, 0.01, 0.05, 0.1, 0.3} on distance metrics in Table 8 and Table 8. For LR, training was performed for 15 epochs with a stepsize of 0.05 and batch size of 32. For CNN, training was carried out for 20 epochs with a step size of 0.05 and a batch size of 64. (ii) Furthermore, in Figure 9, to investigate the effect of epochs under small step sizes, LR was trained for 100 epochs and 200 epochs with a step size of 0.01 (initially 0.05 with 15 epochs). For CNN, training for 50 epochs and 500 epochs with a step size of 0.025 (originally 0.05 with 20 epochs) was performed.
@@
Impact of step sizes. As shown in Table 8 and Table 8, a smaller step size enables our model to be closer to the retrained model. Specifically, as Theorem 4 demonstrates that, in this process, the error arises from the scaled remainder term, i.e., the term $o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})$ is scaled by the subsequent $\mathbf{M}_{e,b(u)}$. Therefore, the approximation error mainly consists of two components: (1) the step size $\eta_{e,b}$ and (2) the maximum eigenvalue $\lambda_{1}$ of $\mathbf{H}_{e,b}$. A smaller step size $\eta$ contracts the eigenvalues of the matrix $\eta_{e,b}\mathbf{H}_{e,b}$, thereby reducing the spectral radius $\rho$ of the matrix $\mathbf{I}-\eta_{e,b}\mathbf{H}_{e,b}$, and ultimately reducing the approximation error.
@@
Table 7: Impact of Step Sizes on Convex Setting. Stepsize $0.05$ represents the hyperparameter choice of experiments for LR in the main paper. It can be observed that as the step size decreases, the unlearned model is closer to the retrained model. (Seed 42)
@@
0.005 0.018 0.040 0.092 0.128
0.01 0.023 0.0518 0.127 0.178
0.05 0.027 0.064 0.149 0.210
0.1 0.028 0.066 0.154 0.217
0.3 5.581 17.48 28.45 40.64
Table 8: Impact of step sizes on non-Convex setting. Stepsize $0.05$ represents the hyperparameter choice of experiments for CNN in the main paper. It can be observed that as the step size decreases, the unlearned model is closer to the retrained model. (Seed 42)
@@
0.1 0.437 0.795 1.722 2.72
0.3 122.2 393.4 1647 2455
Furthermore, when the step size is smaller than a certain threshold, increasing the step size does not lead to significant errors. However, when the step size exceeds this threshold (making $\rho$ not less than or close to 1), the errors will grow exponentially, as demonstrated in Table 8 and Table 8 for step size $\eta\geq 0.3$. In this scenario, our method leads to a complete breakdown of the model, rendering it unusable. However, it is noteworthy that theoretical predictions (Wu et al. (2018)) and empirical validation (Gilmer et al. (2022)) suggested that, across all datasets and models, successful training occurs only when optimization enters a stable region of parameter space where $\lambda_{1}\cdot\eta\leq 2$. Therefore, achieving good optimization results typically does not require a strategy of using aggressive stepsizes.
The connection between learning and unlearning. There is a connection between the sharpness of learning generalization studies and the unlearning method that we propose. Specifically, these studies aim to explore the sharpness of the Hessian of loss function $\mathbf{H}_{e,b}$, and the term ‘sharpness’ refers to the maximum eigenvalue of the Hessian, denoted as $\lambda_{1}$ in some studies, such as Gilmer et al. (2022). A flat loss space is widely recognized as beneficial for gradient descent optimization, leading to improved convergence. Importantly, Gilmer et al. (2022) demonstrated the central role that $\lambda_{1}$ plays in neural network optimization, emphasizing that maintaining a sufficiently small $\lambda_{1}$ during optimization is a necessary condition for successful training (without causing divergence) at large step sizes, as shown in Figure 8. Models that train successfully enter a region where $\lambda_{1}\cdot\eta\leq 2$ mid training (or fluctuate just above this bound). In our paper, we leverage the second-order information from the Hessian at each iteration during the training process to compute the approximators and achieve forgetting. As indicated in Theorem 4, a small $\lambda_{1}$ of the Hessian is beneficial for reducing the approximation error. Therefore, effectively reducing $\lambda_{1}$ during optimization, so that the training enters a flatter region, will not only improve the generalization ability of the learning stage model but also reduce the approximation error in the unlearning stage. As suggested in Gilmer et al. (2022), strategies like stepsize warm-up or initialization strategies for architectures can be employed to stabilize learning and potentially reduce unlearning errors by decreasing $\lambda_{1}$, enabling training at larger step sizes.
Figure 8: The connection between learning and unlearning. We conduct experiment on MNIST and keep the other hyperparameters fixed and adjust the learning rate from 0.001 to 0.31. Under a fixed iteration budget, we have: (1) When the step size is small (less than the threshold $\eta_{1}=0.01$), the training is often insufficient, leading to lower test accuracy. (2) Increasing the step size ensures sufficient training, but when the learning rate becomes too large, overfitting and instability lead to a decrease in test accuracy. We also observe the following phenomena in the unlearning process: (1) When the step size is below a threshold $\eta_{2}=0.25$, the growth of approximation error is extremely slow. (2) However, once the step size exceeds this threshold, the error grows exponentially.
Impact of small step size with more epochs. Finally, as shown in Figure 9, with a small step size (0.005) during model training, we observed a lower approximation error. Even with large epoch sizes, the approximation error remains lower compared to cases with a large step size, such as 0.05. Therefore, selecting an appropriately small step size ensures smaller approximation errors, as these errors typically converge to a smaller value as the number of epochs increases.
@@
Figure 9: Impact of Small Step Size with More Epochs. (a)(b) correspond to LR and CNN with 20% of the data to be forgotten, respectively. The red bar represents the distance metric when using a larger step size; the blue bar represents the distance metric with a small step size during the larger epochs. It can be observed that even when training more epochs with smaller stepsize, the approximation error remains lower compared to training with larger stepsize. (Seed 42)